OpenStudy (anonymous):

a line passes through the point (-3,-4) and its y-intercept is (0,9). what is the equation of the line that is parallel to the first line and passes through the point (3,-7)?

OpenStudy (anonymous):

1st use two point form..that is \[\frac{ y-y1 }{ x-x1 }=\frac{ y2-y1 }{ x2-x1 }\] after finding the eq transform it into the form y=mx+c then compare its slope...and use the point(3,-7) to find the eq of req lines..